OpenStudy (anonymous):

How would you graph f(x)=√4x+4 on a number line?

OpenStudy (allank):

Please clarify: graph on a number line or Cartesian plane?

OpenStudy (anonymous):

A Number Line

OpenStudy (anonymous):

If you mean \[f(x) = \sqrt{4x} + 4\] then \[f(x) = \sqrt{4x} + 4 = 2\sqrt{x} + 4\] and the domain is \[x \ge 0\] so the lowest value can be '4' so on the number line you can include '4' and shade in everything to the right side of the number line (positive solutions)

OpenStudy (anonymous):

Is it greater than or equal to, or just greater than. Like < or the one with the line under meaning equal to.

OpenStudy (anonymous):

You want to include '4', so\[x \ge 4\]